OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 2x + 3y = 4 x - 2y = -5

OpenStudy (anonymous):

i believe it would be -1

OpenStudy (anonymous):

or 0

OpenStudy (anonymous):

2x + 3y = 4 x - 2y = -5 x = -5 + 2y 2(-5 +2y) + 3y = 4 -10 + 4y +3y = 4 -10 +7y = 4 7y = 14 y = 2 2x + 6 = 4 2x = -2 x = -1 x - 4 = -5 x = -1

OpenStudy (anonymous):

multiply 2nd equation with 2 and substract from 1st equation. then you will get 2y=14 y=7. now put value of y in any eqn and solve for x